OpenStudy (anonymous):

Please I really need help. I've already gone to tutoring & they didn't really help... A particle in one dimensional motion starts from rest and reaches a speed of 10 m/sec after 1 sec. During that time its acceleration steadily increases. Is the average velocity during that 1 second, less than, greater than, or equal to 5 m/sec? Explain

OpenStudy (anonymous):



OpenStudy (anonymous):

=5

OpenStudy (anonymous):

umm, yeah that's not it.

OpenStudy (anonymous):

so, I know that the velocity will be less than 5 but I don't know how to explain it. The acceleration is not constant and is going in a linear rate, so the velocity is moving exponentially because of the antiderivative...

OpenStudy (aaronq):

\(v=\dfrac{\Delta x}{\Delta t}=\dfrac{(10-0)m}{(1-0)s}=10m/s\)

OpenStudy (anonymous):

f=1/2mv^2